(12 points) Suppose there are two types of people: high-ability and low-ability. The schooling required to...

  1. (12 points) Suppose there are two types of people: high-ability and low-ability. The schooling required to obtain a particular diploma costs a high-ability person $8,000 per year, and it costs a low-ability person $10,001 per year. Firms wish to use this diploma as a signal of an individual’s ability. The firms intend to pay workers such that the present value of lifetime earnings for a worker without the diploma is $250,000, and the present value of lifetime earnings for a worker with the diploma is $450,000.

  1. What range of years of education must be required to obtain this diploma in order for the diploma to be an effective signal of worker ability?
  2. In equilibrium, in a competitive labor market with workers making rational decisions, how many years of education will be required to obtain this diploma? Briefly explain.
  3. Consider this model relative to the model of learned human capital. Briefly discussion the implications of each model for public funding of education.
Answer #1

Given low ability fee is $10001

Given high ability fee is $8000

Therefore value of skilll $2001

Now given value with diploma is $450000

and value without diploma is $250000

Therefore value of diploma is $200000

Time line for high ability is = 200000 / 8000

= 25 years

Time line for low ability is = 200000 / 10000

= 20 years

Answer (a)  range of years of education must be required to obtain this diploma in order for the diploma to be an effective signal of worker ability is 20-25 years

For equilibrium, in a competitive labor market

No. of years required to equalize the gap would be = Value of diploma / Value of skill

= 200000 / 2001

= 99.95 years or 100 years in round off [Answwer (b)]

Answer (c)

Learned Human capital is the stock of habits, knowledge, social and personality attributes (including creativity) embodied in the ability to perform labour so as to produce economic value.

In assessment of the current and emerging approaches used by state governments in allocating funding for higher education institutions and programs. It reviews a number of desired characteristics or outcomes for state higher education funding models, including equity, adequacy, stability, and flexibility. Although there is certainly no one best approach, given that each state has its own unique set of policy issues and constraints, but current period addresses the types of questions and issues that should be considered in a thoughtful and deliberate manner by state policymakers and higher education leaders in addressing the important, perennial issue of how increasingly limited public funds can be allocated to the greatest benefit for all in the education field. Implications of all these models helps to grow our learned human capital